# BhasaAnuvaad: A Speech Translation Dataset for 14 Indian Languages

### v1

@@ Abstract
Automatic Speech Translation (AST) datasets for Indian languages remain critically scarce, with public resources covering fewer than 10 of the 22 official languages. This scarcity has resulted in AST systems for Indian languages lagging far behind those available for high-resource languages like English. In this paper, we first evaluate the performance of widely-used AST systems on Indian languages, identifying notable performance gaps and challenges. Our findings show that while these systems perform adequately on read speech, they struggle significantly with spontaneous speech, including disfluencies like pauses and hesitations. Additionally, there is a striking absence of systems capable of accurately translating colloquial and informal language, a key aspect of everyday communication. To this end, we introduce BhasaAnuvaad, the largest publicly available dataset for AST involving 14 scheduled Indian languages spanning over 44,400 hours and 17M text segments. BhasaAnuvaad contains data for English speech to Indic text, as well as Indic speech to English text. This dataset comprises three key categories: (1) Curated datasets from existing resources, (2) Large-scale web mining, and (3) Synthetic data generation. By offering this diverse and expansive dataset, we aim to bridge the resource gap and promote advancements in AST for low-resource Indian languages, especially in handling spontaneous and informal speech patterns.
BhasaAnuvaad: A Speech Translation Dataset for 14 Indian Languages
@@ 1 Introduction
Figure 1: Per language statistics of the BhasaAnuvaad dataset for Speech Translation across 14 Indian languages, depicting number of thousands of hours paired with text translations.
Automatic Speech Translation (AST) has become crucial in enabling communication across languages and cultures by breaking down language barriers. Traditionally, speech translation systems have relied on cascaded architectures, where automatic speech recognition (ASR) is followed by machine translation (MT). However, recent advances have led to more integrated end-to-end (E2E) models Babu et al. (2021); Pratap et al. (2024); Radford et al. (2023); Communication et al. (2023) that directly translate speech from one language into text in another. Additionally, Audio-LLM-based systems have emerged Wu et al. (2023); Chu et al. (2023); Fathullah et al. (2023), further advancing the field. These improvements have mainly benefited high-resource languages like English and European languages, where abundant data and well-established benchmarks have driven significant progress.
@@
A critical issue in AST research is the predominant focus on evaluating systems using read speech benchmarks like FLUERS Conneau et al. (2022). While these benchmarks are useful, they fail to capture the complexities of spontaneous, real-world speech. Unlike read speech, spontaneous speech includes informal language, hesitations, extensive code-switching Shankar et al. (2024), and unpredictable structures, all of which are essential for building systems that can handle real-world use cases. Additionally in the case of FLEURS, the speech content has been recorded by only three native speaker, which doesn’t efficiently capture the dialectal variations present in the languages across India. Hence, systems that perform well on read speech often struggle with spontaneous speech, limiting their real-world applicability.
@@
To address this gap, we introduce BhasaAnuvaad, the largest Indic-language AST dataset spanning over 44,400 hours of speech and 17M text segments (Figure 1 and 3). This dataset includes mined parallel speech data, high-quality curated parallel speech data collected from minable corpora, and large-scale synthetic ST data generated using IndicVoices Javed et al. (2024), specifically designed to capture spontaneous speech in Indic languages (Figure 2). Out of the 44,400 hours, around 36,600 hours of data have been created through this project and released into the public domain. This includes about 7,400 hours of data from the SEAMLESS-Align project, where we replicated the intricate mining pipeline to collect the mined data. This diverse dataset provides the resources needed to develop more robust AST systems capable of handling spontaneous speech. In summary, our contributions are: (i) Indic-Spontaneous-Synth, a one of a kind created benchmark for evaluating spontaneous speech translations, (ii) an evaluation of popular AST baselines on both read and spontaneous benchmarks, and (iii) the release of BhasaAnuvaad, the largest Indic-language speech translation corpus, comprising over 44 thousand hours of data across 15 languages. All code and datasets developed as part of this work will be made publicly available to support future research in AST for Indic languages.
@@ 2 Related Works
Our work focuses on speech translation dataset creation for Indian languages. Cascaded AST. The traditional cascaded approach to Automatic Speech Translation (AST) has been a fundamental technique explored extensively in early research Stentiford and Steer (1988); Di Gangi et al. (2019a); Cheng et al. (2019); Bahar et al. (2020). This method connects Automatic Speech Recognition (ASR) and Machine Translation (MT) systems in sequence, which, while effective, is known to suffer from several limitations. Despite its straightforward and simple design, these systems are prone to issues such as cascaded error propagation Sperber and Paulik (2020); Beck et al. (2019), loss of prosodic information Bentivogli et al. (2021), and increased computational complexity Lee et al. (2022).
End-to-End AST. In contrast, end-to-end (E2E) models have gained traction due to their ability to be jointly trained, thereby eliminating the intermediate representations and reducing error propagation Weiss et al. (2017); Inaguma et al. (2021). However, these systems still face challenges, particularly with low-resource languages due to the lack of good quality data Sethiya and Maurya (2023). With the rise of modern Large Language Models (LLMs), recent works have started exploring the feasibility of using Audio-based LLMs (or Speech LLMs), which combine Speech Foundation models (SFMs) and LLMs Fathullah et al. (2023); Wu et al. (2023); Huang et al. (2023).
@@
Datasets for AST. The collection of speech-translation datasets for training end-to-end models has proven to be a persistent challenge. While efforts like SpeechMatrix Duquenne et al. (2022) and SeamlessAlign Communication et al. (2023) attempt to mine data from the web, the resulting datasets are often noisy. Most research has focused on curating data using comparable corpora from domains like Education Ďurišková et al. (2024); Song et al. (2019), TED talks Salesky et al. (2021), Parliament Debates Koehn (2005), among others. Additionally, with the rise of generative models, recent works have also started exploring the generation of large amounts of synthetic speech translation data in a flexible and cost-effective manner Bamfo Odoom et al. (2024). Although the use of synthetic data has been widely studied in text-based machine translation Sennrich et al. (2016); Gala et al. (2023), its application in Speech translation, particularly for low-resource languages like Indic languages, remains underexplored.
@@ 3 Indic-Spontaneous-Synth
Creating a benchmark for spontaneous, real-world speech translation is particularly challenging. Although there have been efforts to develop such datasets—such as Europarl-ST Iranzo-Sánchez et al. (2019), MuST-C Di Gangi et al. (2019b), and mTEDx Salesky et al. (2021)—there are no similar alternatives for Indian languages. To bridge this gap, we investigate the development of a synthetic benchmark utilizing LLMs, as illustrated in Figure 2. While our approach is not intended as a definitive solution, it emphasizes the current lack of resources for Indic languages rather than presenting an ideal benchmark. This method allows us to create a first-of-its-kind benchmark for Indian languages, following a similar approach used in previous synthetic benchmark efforts Lai et al. (2023). We use this as a stepping stone towards addressing the broader lack of resources in this area.
We introduce Indic-Spontaneous-Synth  a comprehensive benchmark covering 14 of the 22 scheduled Indian languages in the XX-to-English (XX-En) direction and 10 Indian languages in the English-to-XX (En-XX) direction. For the XX-En direction, we use IndicVoices Javed et al. (2024), a dataset that features natural, spontaneous speech and captures the cultural, linguistic, and demographic diversity of India. Since the test set of IndicVoices is not publicly available, we create Indic-Spontaneous-Synth using its dev set. We take the “normalized” transcripts from IndicVoices and prompt llama-3.1-405b-instruct model to generate their corresponding clean English translation. We deliberately chose not to use any specialized translation models like IndicTrans2 Gala et al. (2023), as these models are primarily trained for formal text-based translations and lack the capability to handle colloquial speech well. In contrast, with llama-3.1-405b-instruct we obtain steerable translations that perform well in informal and conversational contexts. Finally, we pair the English translations with their corresponding Indic audios, resulting in a dataset with a total of 42.53 hours of audio-text pairs across 14 Indian languages.
@@ 5 BhasaAnuvaad
Total 108.0 62.8 7462.3 3006.0 1076.7 308.6 499.0 298.7 456.4 226.5 6626.5 2603.3 3988.2 288.0
@@
To overcome the challenges with spontaneous speech, we introduce BhasaAnuvaad , a large-scale, high-quality dataset covering 14 of the 22 scheduled Indian languages across a wide array of domains, use cases and styles.
@@ 5.2.2 Sources of Mining Corpora
Mann ki Baat: Mann Ki Baat [^15] https://www.pmindia.gov.in/ is a weekly radio podcast organized by the Indian Prime Minister to discuss various national Issues. The podcast is transcribed and manually translated into 13 Indian languages. It is then recorded in these languages, also through manual efforts, resulting in over 425 hours of speech data for both Indic-En and En-Indic directions. Notably, Mann Ki Baat is the only source in this dataset that includes Manipuri, though it is presented in Bengali script.
@@ 5.3 Synthetic Data Generation
To ensure that we have audio diversity across different domains, demographics, and languages, we use IndicVoices which is the largest collection of natural and spontaneous speech containing over 4000 hours of transcribed speech. Its extensive language diversity makes it well-suited for training models capable of performing effectively in real-world scenarios.
@@
Building on the insights gained from evaluating existing systems and recognizing their shortcomings, we introduced BhasaAnuvaad —a comprehensive, large-scale dataset designed to bridge the performance gap in real-world spoken translation applications. With BhasaAnuvaad, we aim to address the complexities of spontaneous speech across the 14 of the 22 scheduled languages. By incorporating diverse domains and varied speaking styles, BhasaAnuvaad  aims to support the development of models capable of meeting the dynamic requirements of deployment environments. Through this resource, we aim to not only enhance the robustness of Spoken Translation systems but also provide a foundational tool that supports future advancements in multilingual speech translation for Indian languages.
@@ 6 Conclusion
In this work, we evaluated popular Spoken Translation systems for Indian languages and identified key limitations in their ability to handle spontaneous speech, a common occurrence in real-world deployment scenarios. Our findings underscore the need for more rigorous evaluation and benchmarking of these systems in practical settings. To address this gap, we introduced a novel benchmark,  Indic-Spontaneous-Synth, comprising 130 hours of spontaneous speech data across 10 languages, covering diverse domains and tasks to provide a more realistic testbed for system performance. Additionally, recognizing the scarcity of resources covering multiple languages, domains, and speaking styles, we presented  BhasaAnuvaad-—a large-scale, meticulously curated dataset featuring over 44,000 hours of speech data and 17 million text segments across 10 languages. This dataset includes a broad spectrum of practically relevant domains and styles, offering a robust foundation for training models.

### v2

# BhasaAnuvaad: A Speech Translation Dataset for 13 Indian Languages

@@ Abstract
Automatic Speech Translation (AST) datasets for Indian languages remain critically scarce, with public resources covering fewer than 10 of the 22 official languages. This scarcity has resulted in AST systems for Indian languages lagging far behind those available for high-resource languages like English. In this paper, we first evaluate the performance of widely-used AST systems on Indian languages, identifying notable performance gaps and challenges. Our findings show that while these systems perform adequately on read speech, they struggle significantly with spontaneous speech, including disfluencies like pauses and hesitations. Additionally, there is a striking absence of systems capable of accurately translating colloquial and informal language, a key aspect of everyday communication. To this end, we introduce BhasaAnuvaad, the largest publicly available dataset for AST involving 13 out of 22 scheduled Indian languages and English spanning over 44,400 hours and 17M text segments. BhasaAnuvaad contains data for English speech to Indic text, as well as Indic speech to English text. This dataset comprises three key categories: (1) Curated datasets from existing resources, (2) Large-scale web mining, and (3) Synthetic data generation. By offering this diverse and expansive dataset, we aim to bridge the resource gap and promote advancements in AST for Indian languages.
BhasaAnuvaad: A Speech Translation Dataset for 13 Indian Languages
@@ 1 Introduction
Figure 1: Statistics of the BhasaAnuvaad dataset for Speech Translation across 13 Indian languages, depicting number of audio hours paired with text translations.
Automatic Speech Translation (AST) has become crucial in enabling communication across languages and cultures by breaking down language barriers. Traditionally, speech translation systems have relied on cascaded architectures, where Automatic Speech Recognition (ASR) is followed by Machine Translation (MT). However, recent advances have led to more integrated end-to-end (E2E) models Babu et al. (2021); Pratap et al. (2024); Radford et al. (2023); Communication et al. (2023) that directly translate speech from one language into text in another. Additionally, Audio-LLM-based systems have emerged Wu et al. (2023); Chu et al. (2023); Fathullah et al. (2023); Gaido et al. (2024) further advancing the field. These improvements have mainly benefited high-resource languages like English and European languages, where abundant data and well-established benchmarks have driven significant progress.
@@
A critical issue in AST research is the predominant focus on evaluating systems using read speech benchmarks like FLUERS Conneau et al. (2022). While these benchmarks are useful, they fail to capture the complexities of spontaneous, real-world speech. Unlike read speech, spontaneous speech includes informal language, hesitations, extensive code-switching Shankar et al. (2024), and unpredictable structures, all of which are essential for building systems that can handle real-world use cases. Additionally, in the case of FLEURS, the speech content has been recorded by only three native speakers, which doesn’t efficiently capture the dialectal variations present in the languages across India. Hence, systems that perform well on read speech often struggle with spontaneous speech, limiting their real-world applicability.
@@
To address this gap, we introduce BhasaAnuvaad, the largest Indic-language AST dataset spanning over 44,400 hours of speech and 17M text segments (Figure 1 and 3) for 13 of 22 scheduled Indian languages and English. The 13 Indian languages are: Assamese, Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Nepali, Odia, Punjabi, Tamil, Telugu, and Urdu. This dataset includes mined parallel speech data, high-quality curated parallel speech data collected from existing corpora, and large-scale synthetic ST data generated using IndicVoices Javed et al. (2024), specifically designed to capture spontaneous speech in Indic languages (Figure 2). Out of the 44,400 hours, around 36,600 hours of data have been newly created or collected and released into the public domain. This includes about 7,400 hours of data from the SEAMLESS-Align project, where we replicated the intricate mining pipeline to collect the mined data. This diverse dataset provides the resources needed to develop more robust AST systems capable of handling spontaneous speech.
In summary, our contributions are: (i) Indic-Spontaneous-Synth, a one of a kind created benchmark for evaluating spontaneous speech translations, (ii) an evaluation of popular AST baselines on both read and spontaneous benchmarks, and (iii) the release of BhasaAnuvaad, the largest Indic-language speech translation corpus, comprising over 44 thousand hours of data across 13 Indian languages and English. All code and datasets developed as part of this work will be made publicly available to support future research in AST for Indic languages.
@@ 2 Related Works
Our work focuses on speech translation dataset creation for Indian languages.
Cascaded AST. The traditional cascaded approach to Automatic Speech Translation (AST) has been a fundamental technique explored extensively in early research Stentiford and Steer (1988); Di Gangi et al. (2019a); Cheng et al. (2019); Bahar et al. (2020). This method connects Automatic Speech Recognition (ASR) and Machine Translation (MT) systems in sequence, which, while effective, is known to suffer from several limitations. Despite its straightforward and simple design, these systems are prone to issues such as cascaded error propagation Sperber and Paulik (2020); Beck et al. (2019), loss of prosodic information Bentivogli et al. (2021), and increased computational complexity Lee et al. (2022).
End-to-End AST. In contrast, end-to-end (E2E) models have gained traction due to their ability to be jointly trained, thereby eliminating the intermediate representations and reducing error propagation Weiss et al. (2017); Inaguma et al. (2021). However, these systems still face challenges, particularly with low-resource languages due to the lack of good quality data Sethiya and Maurya (2023). With the rise of modern Large Language Models (LLMs), recent works have started exploring the feasibility of using Audio-based LLMs (or Speech LLMs), which combine Speech Foundation models (SFMs) and LLMs Fathullah et al. (2023); Wu et al. (2023); Huang et al. (2023); Gaido et al. (2024) .
@@
Datasets for AST. The collection of speech-translation datasets for training end-to-end models has proven to be a persistent challenge. While efforts like SpeechMatrix Duquenne et al. (2022) and SeamlessAlign Communication et al. (2023) attempt to mine data from the web, the resulting datasets are often noisy. Most research has focused on curating data using comparable corpora from domains like Education Ďurišková et al. (2024); Song et al. (2019), TED talks Salesky et al. (2021), Parliament Debates Koehn (2005), among others. Additionally, with the rise of generative models, recent works have also started exploring the generation of large amounts of synthetic speech translation data in a flexible and cost-effective manner Ye et al. (2023)Bamfo Odoom et al. (2024). Although the use of synthetic data has been widely studied in text-based machine translation Sennrich et al. (2016); Gala et al. (2023), its application in Speech translation, particularly for low-resource languages like Indic languages, remains underexplored.
@@ 3 Indic-Spontaneous-Synth
Creating a benchmark for spontaneous, real-world speech translation is particularly challenging. Although there have been efforts to develop such datasets—such as GigaST Ye et al. (2023), Europarl-ST Iranzo-Sánchez et al. (2019), MuST-C Di Gangi et al. (2019b), and mTEDx Salesky et al. (2021)—there are no similar alternatives for Indian languages. To bridge this gap, we investigate the development of a synthetic benchmark utilizing LLMs, as illustrated in Figure 2. While our approach is not intended as a definitive solution, it emphasizes the current lack of resources for Indic languages rather than presenting an ideal benchmark. This method allows us to create a first-of-its-kind benchmark for Indian languages, following a similar approach used in previous synthetic benchmark efforts Lai et al. (2023). We use this as a stepping stone towards addressing the broader lack of resources in this area.
We introduce Indic-Spontaneous-Synth  a comprehensive benchmark covering 13 of the 22 scheduled Indian languages in the XX-to-English (XX-En) direction and 10 Indian languages in the English-to-XX (En-XX) direction. For the XX-En direction, we use IndicVoices Javed et al. (2024), a dataset that features natural, spontaneous speech and captures the cultural, linguistic, and demographic diversity of India. Since the test set of IndicVoices is not publicly available, we create Indic-Spontaneous-Synth using its dev set. We take the “normalized” transcripts from IndicVoices and prompt llama-3.1-405b-instruct model to generate their corresponding clean English translation. We deliberately chose not to use any specialized translation models like IndicTrans2 Gala et al. (2023), as these models are primarily trained for formal text-based translations and lack the capability to handle colloquial speech well. In contrast, with llama-3.1-405b-instruct we obtain steerable translations that perform well in informal and conversational contexts. Finally, we pair the English translations with their corresponding Indic audios, resulting in a dataset with a total of 42.53 hours of audio-text pairs across 13 Indian languages, namely Assamese (asm), Bengali (ben), Gujarati (guj), Hindi (hin), Kannada (kan), Malayalam (mal), Marathi (mar), Nepali (npi), Odia (ory), Punjabi (pan), Tamil (tam), Telugu (tel) and Urdu (urd).
@@ 5 BhasaAnuvaad
Total 108.0 90.0  [^8] The Khan Academy Corpus, although published is yet to be released publicly. We will update the statistics once the data is accessible to the public. 7462.3 3006.0 1076.7 308.6 499.0 298.7 456.4 226.5 6626.5 2603.3 3988.2 288.0
@@
To overcome the challenges with spontaneous speech, we introduce BhasaAnuvaad , a large-scale, high-quality dataset covering 13 of the 22 scheduled Indian languages and English across a wide array of domains, use cases and styles.
@@ 5.2.2 Sources of Mining Corpora
Mann ki Baat: Mann Ki Baat [^15] https://www.pmindia.gov.in/ is a weekly radio podcast organized by the Indian Prime Minister to discuss various national Issues. The podcast is transcribed and manually translated into 13 Indian languages. It is then recorded in these languages, also through manual efforts, resulting in over 425 hours of speech data for both Indic-En and En-Indic directions. Notably, Mann Ki Baat is the only source in this dataset that includes Manipuri, though it is presented in Bengali script. We note that, although the tables and figures do not specifically present it, we are releasing Manipuri in Bengali script too as sourced from Mann Ki Baat.
@@ 5.3 Synthetic Data Generation
To ensure that we have audio diversity across different domains, demographics, and languages, we use IndicVoices which is the largest collection of natural and spontaneous speech containing over 4,000 hours of transcribed speech. Its extensive language diversity makes it well-suited for training models capable of performing effectively in real-world scenarios.
@@
Building on the insights gained from evaluating existing systems and recognizing their shortcomings, we introduced BhasaAnuvaad —a comprehensive, large-scale dataset designed to bridge the performance gap in real-world spoken translation applications. With BhasaAnuvaad, we aim to address the complexities of spontaneous speech across 13 of the 22 scheduled languages. By incorporating diverse domains and varied speaking styles, BhasaAnuvaad  aims to support the development of models capable of meeting the dynamic requirements of deployment environments. Through this resource, we aim to not only enhance the robustness of Spoken Translation systems but also provide a foundational tool that supports future advancements in multilingual speech translation for Indian languages.
@@ 6 Conclusion
In this work, we evaluated popular Spoken Translation systems for Indian languages and identified key limitations in their ability to handle spontaneous speech, a common occurrence in real-world deployment scenarios. Our findings underscore the need for more rigorous evaluation and benchmarking of these systems in practical settings. To address this gap, we introduced a novel benchmark,  Indic-Spontaneous-Synth, comprising 120 hours of spontaneous speech data across 13 Indian languages and English, covering diverse domains and tasks to provide a more realistic testbed for system performance. Additionally, recognizing the scarcity of resources covering multiple languages, domains, and speaking styles, we presented  BhasaAnuvaad-—a large-scale, meticulously curated dataset featuring over 44,400 hours of speech data and 17 million text segments across 13 Indian languages and English. This dataset includes a broad spectrum of practically relevant domains and styles, offering a robust foundation for training models.